The base of a triangle is 4cm greater than the height. the area is 48cm^2. Find the height and length of the base!

Thank you

Answer by checkley77(12844)

B=H+4
A=BH/2
48(H+4)H/2
48=(H^2+4H)/2
H^2+4H=48*2
H^2+4H=96
H^2+4H-96=0
(H+12)(H-8)=0
H-8=0
H=8 ANSWER FOR THE HEIGHT.
B=8+4=12 ANS. FOR THE HEIGHT.
PROOF:
48=8*12/2
48=96/2
48=48